# The Shape of Word Embeddings: Recognizing Language Phylogenies through Topological Data Analysis

### v1

@@ 3.2 Constructing Algorithmic Phylogenies
The NJ method Saitou and Nei (1987) defines distances between clusters, $D(\cdot,\cdot)$, inductively through a merging process. Let $C$ be the collection of current clusters. For each cluster, $c\in C$, define
$\alpha(c)\coloneqq\frac{1}{|\,{C}\,|-2}\sum_{\begin{subarray}{c}d\in C\\ d\neq c\end{subarray}}D(c,d).$
Find a pair $a,b\in C$ minimizing $D(a,b)-\alpha(a)-\alpha(b)$. We replace $a,b$ by the union, $a\cup b$, and define the new distances $D(a\cup b,d)\coloneqq(D(a,d)+D(b,d)-D(a,b))/2$ for each other $d$.
Figure 2 gives examples of reconstructed and reference trees.
(a) Reference Ethnologue tree
(b) UPGMA
(c) NJ
@@ 4 Analysis Pipeline
We work with language embeddings from FastText Grave et al. (2018). We filter the data in two ways. Firstly, we use the $V=10,000$ most frequent tokens of each language. Secondly, we only choose a subset of the languages to work with. Because low-resource languages may skew our analysis, we filter the languages by the number of Wikipedia articles [^2] https://meta.wikimedia.org/wiki/List_of_Wikipedias, extracted on November 8, 2023, and consider the first 30, 50 or all 81 Indo-European languages.
Embedding distance matrices.
The persistent homology pipeline we use takes a distance matrix as the input. Therefore, for each language, we generate a $V\times V$ matrix, each entry being the distance between two 300-dimensional vectors, $u$, $v$. We use two different notions of distances—this is the second parameter—either the standard Euclidean distance, $||\,{u-v}\,||$, or the cosine distance, $1-a\cdot b\,/\left(||\,{a}\,||\cdot||\,{b}\,||\right)$, which is often used in language processing.
@@
For the embedding distance matrix of each language, we compute the persistence diagram of its Vietoris-Rips complex, as described in Section 2.1, using an efficient implementation of this computation, Ripser Bauer (2021) [^3] https://github.com/Ripser/ripser. In this step, we choose the dimension of the topological features to use—either 0, 1 or 2.
@@
We use bottleneck distance implementation from GUDHI library [^4] https://gudhi.inria.fr François (2016), and slice Wasserstein and persistence image implementations from persim Python library [^5] https://github.com/scikit-tda/persim. Bar statistics are computed directly.
For the bars statistics, in dimension 00 we only use the death, $d$, value, as $b=0$ for all dots. That is, the vectorization is 10-dimensional in this case rather than 40-dimensional as in the other cases.
@@
Finally, we construct a phylogenetic tree based on the language distance matrix. We try two different approaches. First, we construct the tree with either of the two agglomerative clustering algorithms described in Section 3—neighbor joining (NJ) or unweighted pair group method with arithmetic mean (UPGMA). Both of those algorithms are implemented in biopython package [^6] https://biopython.org/wiki/Phylo under DistanceTreeConstructor class.
We also conducted a second set of experiments to allay concerns about the difference in height/topology between algorithm and reference trees, by fixing our algorithmic trees to have the same unlabeled topology as the reference tree. The random labeling whose tree distances between language pairs best correlated with the TDA-based distances between language word embeddings is taken as our algorithmically-designed tree, which can then be compared to the Ethnologue tree using the distance metrics previously defined. Consistent with the results to be reported below, we perform substantially better than chance for almost all combinations of TDA variant and tree distance metric, typically by two or more standard deviations from the mean. Details are presented in the Appendix.
Figure 3: The statistical significance of TDA trees for 30, 50, and 81 languages against the Ethnologue reference for trees reconstructed by UPGMA and NJ for each combination of parameters described in Section 4. Each dot represents the distance of a single reconstructed tree to the reference, and its position shows how many standard deviations from the mean it lies in the distance distribution over 100,000 random permutations of the tree’s leaves.
5 Experiments
@@ 5.1 Evaluating Phylogenetic Trees
To compare the reference Ethnologue tree $R$ to each phylogenetic tree $T$ constructed using language distances inferred by TDA, we employ six different tree distances, using implementations from the R package TreeDist version 2.6.3. These include (i) path distance Farris (1969); Steel and Penny (1993), (ii/iii) Jaccard-Robinson-Foulds distance for $k=1$ and $k=2$ Nye et al. (2005); Böcker et al. (2013), (iv) matching split distance Bogdanowicz and Giaro (2012); Lin et al. (2012), and (v/vi) the phylogenetic and clustering information distances Smith (2020). These distances are described formally in Appendix Section A.2.
@@
To assess whether the reconstructed trees capture part of the real phylogeny, we evaluate them in terms of leaf label permutations—comparing the distance $d(T,R)$ between the algorithmic tree $T$ and reference $R$ to the distribution of $d(T^{\prime},R)$, where $T^{\prime}$ is obtained from $T$ by shuffling the leaf labels.
@@
For each tree, $T$, constructed by UPGMA or NJ algorithm, we performed 100,000 permutations of its leaf labels. For each permuted tree, we measure the six tree distances to the reference Ethnologue tree with the corresponding number of leaves—30, 50 or 81. Then, separately for each of the six tree distances, we identify where in this distribution of distances $d(T,R)$ lies. The measure of success is to check how many random permutations did worse than $T$. Those numbers are clustered towards the upper end—for 484 out of 864 distributions, at least 95,000 (95%) of the permutations are further from $R$ than $T$ is.
5.2 Results
@@
Figure 3 is the primary result in this paper, presenting the results of evaluations of 144 different algorithmic tree reconstructions using TDA, each evaluated on six different tree similarity measures, for a total of 864 different points. By employing permutation tests, we can map each such point to a significance level, expressed in terms of (a) the fraction of random samples dominated by the reconstructed tree or (b) the number of standard deviations the reconstructed tree sits from the mean value over 100,000 random trees; the options are tightly correlated. Figure 3 summarises (b); analogous figure for (a) is in Appendix Figure 8.
Figure 4: The fraction of parameter combinations (out of 288) in Figure 3 that are bested by at most $p\cdot 100,\!000$ permutations.
@@
Further summary of these results is in Figure 4. The cumulative distribution plots show the fraction of experimental conditions which satisfy a given permutation test $p$-value. With respect to this value, 484 out of the 864 conditions proved significant to at least the $0.05$-level. Of these, 255 were significant to at least $0.01$, 57 to at least $0.001$, and 10 to $0.0001$. In one case the algorithmic tree dominated all $t=100,\!000$ random permutations evaluated. Our statistically strongest results came when using $n=50$ languages: large enough to encode topologies hard to find by chance while avoiding low resource languages with poorly shaped embeddings. Twice as many tests for UPGMA trees (114) were significant to the $0.005$-level as NJ trees (63), demonstrating that better combinatorial optimization leads to more significant phylogenies. Detailed breakdowns are provided in Tables 1, 2, 3, 4 in the Appendix.
@@
That our best tree beats all trial $t$ permutations in our experiment implies it is significant beyond the $1/t$ level. By measuring the quality score of each tree with respect to each distance in terms of the number of standard deviations above its random background mean and assuming normality, we can obtain a $p$-value associated with each such $z$-score. The maximum achieved $6.87\sigma$ corresponds to a Bonferroni-corrected $p$-value of $2.77\times 10^{-9}$. Coupled with all other presented permutation test results, clearly TDA is picking up real signal from the shape of word embeddings.
@@
However, there are reasons to be optimistic that the strength of this signal may be increased with further research and experience. We note that our TDA analysis is “blind”, with no statistical normalizations of numerical range across the different language embeddings. Orthogonal transformations preserve distances and angles, so they do not influence persistent homology, but scaling does change Euclidean distances, and translation can change the cosine distances.
@@
There is somewhat of a gap between the low-dimensional persistent homologies of dimensions $0\leq d\leq 2$ which are computationally accessible and the 300-dimensional word vectors we analyze. In hopes of capturing more of the higher order behavior of high-dimensional data, data is often first reduced to fewer dimensions before persistent homology is applied. Standard methods like PCA, t-SNE or U-Map can be used to reduce dimension to $\sim 10$ or $\sim 50$.
@@
Our results demonstrate that TDA invariants capture properties of natural languages, but it remains completely open about what aspects of language they are keying on. What do tunnels and voids in embedding space tell us about languages? An interesting possible future work revolves around optimizing the parameters to compare persistence diagrams for informative regions to try to identify which structures have concrete linguistic meaning.
@@ Appendix A Supplemental Material
A.1 Preserving the Ethnologue Tree Topology
@@
Thus we investigated an optimization procedure to construct TDA-based trees consistent with the reference tree topology. Instead of constructing a new tree, we try to find the best assignment of labels (languages) to the leaves based on the given distance matrix. A tree always yields a distance matrix between the leaves—the distance of two leaves is the length of the unique path connecting them. We set up an optimization process to find the label assignment leading to the tree distance matrix that best correlates with our given distance matrix. To compare distance matrices, we use the Pearson correlation coefficient. Therefore, if $E$ is the tree distance matrix of the Ethnologue tree, and $D$ is the distance matrix computing from the pipeline as described above, we want to find a permutation matrix $P$ that maximizes the Pearson correlation coefficient between $P^{T}EP$ and $D$.
@@
A.2 Tree Similarity Metrics

### v2

@@ 3 Constructing Phylogenetic Trees
The NJ method Saitou and Nei (1987) defines distances between clusters, $D(\cdot,\cdot)$, inductively through a merging process. Let $C$ be the collection of current clusters. For each cluster, $a\in C$, define
$\alpha(a)\coloneqq\frac{1}{|\,{C}\,|-2}\sum_{\begin{subarray}{c}b\in C\\ b\neq a\end{subarray}}D(a,b).$
In each step, find a pair $a,b\in C$ minimizing $D(a,b)-\alpha(a)-\alpha(b)$. Then replace $a,b$ by the union, $a\cup b$, and define the new distances $D(a\cup b,c)\coloneqq(D(a,c)+D(b,c)-D(a,b))/2$ for each other $c$. The length of the new tree edge, $(a\cup b,a)$ is defined as $D(a,b)+\alpha(a)-\alpha(b)$, and symmetrically for $(a\cup b,b)$.
@@ 4 Analysis Pipeline
This section details the data, the process of going from the data to trees, and finally the evaluation of the reconstructed trees. Our code for the analysis is published on GitHub. [^4] https://github.com/OnDraganov/shape-of-word-embeddings
4.1 Data
The historical origin of languages has been extensively studied, and much is known, but debate remains vigorous even for the Indo-European languages we study here Gray and Atkinson (2003); Pereltsvaig and Lewis (2015); Longobardi et al. (2013). The reference tree we use comes from Ethnologue [^5] https://www.ethnologue.com, which is the most widely consulted inventory of the world’s languages Lewis et al. (2024). First published in 1951 and now in its 26th edition, Ethnologue currently records data about 7,168 living languages. Although there is no universal agreement on language origins (see e.g. Hammarström (2015) for a critique of Ethnologue), it provides a broadly acceptable reference tree for our evaluation purposes, particularly for the major Indo-European languages we consider here. One downside of this reference is the lack of weights on the edges of the tree—we only get the topology. An alternative for a follow-up research could be to compare to other computationally reconstructed trees that are considered reliable, e.g., Serva and Petroni (2008).
FastText [^6] https://fasttext.cc/ Grave et al. (2018) provides a set of pre-trained, 300-dimensional word embeddings for 157 languages, trained on Common Crawl and Wikipedia. There are 81 Indo-European languages that appear in the Ethnologue reference tree and also have pre-computed FastText embeddings. To protect against the risk that low-resource languages might have less reliable embeddings, we will also evaluate filtrations to the most popular 30 and 50 languages. The Ethnologue reference trees for all filtrations are provided for inspection in Appendix Figure 7.
4.2 From Language Embeddings to a Tree
@@
As described in Section 4.1, we work with language embeddings from FastText. We filter the data in two ways. Firstly, we use the $V=10,\!000$ most frequent tokens of each language—a threshold chosen to be as high as possible considering computational feasibility [^7] Degree 2 PH is expensive; for 00 and $1$, $V$ can be increased.. Secondly, we only choose a subset of the languages to work with. Because low-resource languages may skew our analysis, we filter the languages by the number of Wikipedia articles [^8] https://meta.wikimedia.org/wiki/List_of_Wikipedias, extracted on November 8, 2023, and consider the first 30, 50 or all 81 available Indo-European languages.
Token-to-token distance matrices.
As input to the persistent homology pipeline we use, for each language, we generate a $V\times V$ matrix, each entry being the distance between two 300-dimensional vectors, $u$, $v$. We use two different notions of distances: either the standard Euclidean distance, $||\,{u-v}\,||$, or the cosine distance, $1-a\cdot b\,/\left(||\,{a}\,||\cdot||\,{b}\,||\right)$, which is often used in language processing.
@@
For the token-to-token distance matrix of each language, we compute the persistence diagram [^9] We use Vietoris-Rips complex, described in Section A.1. as described in Section 2, using an efficient implementation of this computation, Ripser [^10] https://github.com/Ripser/ripser (Bauer, 2021). In this step, we choose the degree of the topological features to use—either 0, 1 or 2.
@@
We use bottleneck distance implementation from GUDHI library [^11] https://gudhi.inria.fr (François, 2016), and slice Wasserstein and persistence image implementations from persim Python library [^12] https://github.com/scikit-tda/persim. Bar statistics are computed directly.
Persistence images take several parameters. Our choices depend on the embedding metric used and persistence diagram degree considered. For degrees $1$ and $2$ we use a $10\times 10$ grid. For degree 00 it is just $1\times 10$ grid. The range of radii covered by the grid is $[0,10]$ for Euclidean embedding metric, and $[0,1]$ for cosine. Note that if there is a dot outside of this range in the persistence diagram, the Gaussian is still put over it, and spills into the considered range, so even those dots are still considered in the distance. The weight function is linear in all cases, $w(d-b)=d-b$. Careful optimization of these parameters might lead to better results.
For the bars statistics, in degree 00 we only use the death, $d$, value, as $b=0$ for all dots. That is, the vectorization is 10-dimensional in this case rather than 40-dimensional as in the other cases.
@@
Finally, we construct a phylogenetic tree based on the language distance matrix. We try two different approaches. We use either of the two agglomerative clustering algorithms described in Section 3—neighbor joining (NJ) or unweighted pair group method with arithmetic mean (UPGMA). Both of those algorithms are implemented in biopython package [^13] https://biopython.org/wiki/Phylo under DistanceTreeConstructor class. Figure 8 in Appendix gives examples of reconstructed and reference trees.
Two aspects of this approach make comparisons of thus reconstructed trees to the reference difficult. Firstly, the reference Ethnologue tree has no weights on edges, so we cannot meaningfully use the weights of the reconstructed trees either. Secondly, the Ethnologue tree is flat and neither binary nor rooted, while the reconstructed trees are generally somewhat deep binary trees. Our solution to this difference in topology is to test the reconstructed trees against their copies with permuted leaf labels.
We also conducted a second set of experiments to allay the concerns about the difference in height and topology between reconstructed and reference trees. We fix the topology of the reference Ethnologue tree, and try to assign languages to the leaves—to do that, we optimise the similarity of the path-distance matrix given by the labeled tree to the distance matrix computed using TDA. We then evaluate this optimized labeling against 100,000 random labelings. Consistent with the results reported below, we perform substantially better than chance for almost all combinations of parameters, typically by two or more standard deviations from the mean. Details of this alternative analysis are presented in the Appendix (A.2).
Summary of the parameters.
To construct one tree we choose a number of languages (30, 50, 81), an embedding (Euclidean, cosine), a degree of persistence diagrams (0, 1, 2), a distance between persistence diagrams (bottleneck, sliced Wasserstein, persistence image, bars statistics), and a construction of a tree (UPGMA, NJ)—altogether $3\cdot 2\cdot 3\cdot 4\cdot 2=144$ variants. For each there are six tree distances to evaluate how well we reconstructed the phylogeny, as described below.
Figure 2: The statistical significance of TDA trees for 30, 50, and 81 languages against the Ethnologue reference, $E$, for trees reconstructed by UPGMA and NJ for each combination of parameters described in Section 4. Each dot represents a single reconstructed tree, $T$, and a tree distance, $D$. We performed 100,000 random permutations of the leaves of $T$, and compared each to the reference $E$ using the distance $D$. This yields a distribution with mean $\mu$ and standard deviation $\sigma$. To evaluate the reconstruction $T$, we plot $(\mu-D(T,E))/\sigma$. The higher the value, the better the reconstruction. A star inside a dot signifies that $D(T,E)$ is smaller than 95,500 of the permuted tree distances.
@@ 4.3 Evaluating Phylogenetic Trees
To compare the reference Ethnologue tree $E$ to each phylogenetic tree $T$ constructed using the language distances inferred by TDA, we employ six different tree distances, using implementations from the R package TreeDist version 2.6.3. These include (i/ii) Jaccard-Robinson-Foulds distance for $k=1$ and $k=2$ Nye et al. (2005); Böcker et al. (2013), (iii) matching split distance Bogdanowicz and Giaro (2012); Lin et al. (2012), (iv/v) the phylogenetic and clustering information distances Smith (2020), and (vi) path distance Farris (1969); Steel and Penny (1993). We provide detailed descriptions of these distances in Appendix Section A.3.
@@
To assess whether the reconstructed trees capture part of the real phylogeny, we evaluate them in terms of leaf label permutations—comparing the distance $d(T,E)$ between the algorithmic tree $T$ and reference $E$ to the distribution of $d(T^{\prime},E)$, where $T^{\prime}$ is obtained from $T$ by shuffling the leaf labels.
@@
For each of the 144 constructed trees, $T$, we performed 100,000 permutations of its leaf labels. For each permuted tree, we measure the six tree distances to the reference Ethnologue tree with the corresponding number of leaves—30, 50 or 81. Then, separately for each of the six tree distances, we identify where in this distribution of distances $d(T,E)$ lies. The measure of success is to check how many random permutations did worse than $T$.
5 Results of the Experiments
The experiments show correlation between the word embedding dissimilarity and language phylogeny—for 484 out of 864 distributions, at least 95,000 (95%) of the permutations are further from the reference tree $E$ than the reconstructed tree $T$ is.
@@ 5 Results of the Experiments
Figure 2 is the primary result in this paper, presenting the results of evaluations of 144 different tree reconstructions based on TDA, each evaluated on six different tree similarity measures, for a total of 864 different points. By employing permutation tests, we can map each such point to a significance level, expressed in terms of (a) the fraction of random samples dominated by the reconstructed tree or (b) the number of standard deviations the reconstructed tree sits from the mean value over 100,000 permutations of leaves; the options are tightly correlated. Figure 2 summarises (b); analogous figure for (a) is in Appendix Figure 6.
Figure 3: The fraction of parameter combinations (out of 288) in Figure 2 bested by $\leq p\cdot 100,\!000$ permutations.
@@
Further summary of these results is in Figure 3. The cumulative distribution plots show the fraction of experimental conditions which satisfy a given permutation test $p$-value. With respect to this value, 484 out of the 864 conditions proved significant to at least the $0.05$-level. Of these, 255 were significant to at least $0.01$, 57 to at least $0.001$, and 10 to $0.0001$. In one case the algorithmic tree dominated all $t=100,\!000$ random permutations evaluated.
Our statistically strongest results came when using $50$ languages: possibly large enough to encode topologies hard to find by chance while avoiding low resource languages with poorly shaped embeddings. Twice as many tests for UPGMA trees (114) were significant to the $0.005$-level as NJ trees (63), demonstrating that better combinatorial optimization leads to more significant phylogenies.
Different distances between persistent diagrams seem to fit well with different degree and token-to-token metric. Overall, bars statistics and sliced Wasserstein seem to perform better, while bottleneck is the least successful—this might be explained by its strong sensitivity to outliers, and suggests that the other notions might be better suited for analyses of word embeddings’ (dis)similarities. Detailed breakdowns of the success of different parameters are provided in Tables 1, 2, 3, 4 in the Appendix.
@@
That our best tree beats all trial $t$ permutations in our experiment implies it is significant beyond the $1/t$ level. By measuring the quality score of each tree with respect to each distance in terms of the number of standard deviations above its random background mean and assuming normality, we can obtain a $p$-value associated with each such $z$-score. The maximum achieved $6.87\sigma$ corresponds to a Bonferroni-corrected $p$-value of $8642.77\times 10^{-9}=P\left[x\geq 6.87\right]*$ 864, assuming a normal distribution. Coupled with all other presented permutation test results, clearly TDA is picking up real signal from the shape of word embeddings.
6 Conclusions
Our analysis of FastText word embeddings of 81 Indo-European languages suggests that deviations from pair-wise isometry correlates with the historical origin of the languages—an interesting contribution to the debate about the extent to which the variance between monolingual embeddings is explained by essential differences between the languages versus scarcity of resources and under-training when constructing them.
A natural question is to what extent the presented results generalize. Would we get comparably significant results for word embeddings trained with different algorithms, with different parameters or on a different corpus? And if more careful training should lead to isometric embeddings, as suggested by Vulić et al. (2020), is there an explanation for a correlation of lower quality of embeddings with language phylogeny?
The results suggest that TDA methods can help us better understand the structure of word embeddings, and further research can lead to useful insights. Different parameters can be better suited for comparing word embeddings than the previously used bottleneck lower bound of Gromov-Hausdorff distance. Furthermore, better understanding of the particular detected topological features could lead to interesting insights. For example, looking at a single language embedding, we can look back at concrete loops in degree 1 persistence diagrams and ask whether the collection of words it passes through has some meaningful linguistic interpretation. Or, identifying a large loop in one language, we can ask whether the translation of the words also forms a large loop in other languages.
@@ 6 Conclusions
However, there are reasons to be optimistic that the strength of this signal may be increased with further research and experience. We note that our TDA analysis is “blind”, with no statistical normalizations of numerical range across the different language embeddings. Orthogonal transformations preserve distances and angles, so they do not influence persistent homology, but scaling does change Euclidean distances, and translation (shifts) can change the cosine distances.
@@
There is somewhat of a gap between the low-dimensional persistent homologies of dimensions $0\leq d\leq 2$ which are computationally accessible and the 300-dimensional word vectors we analyze. In hopes of capturing more of the higher order behavior of high-dimensional data, data is often first reduced to fewer dimensions before persistent homology is applied. Standard methods like PCA, t-SNE or U-Map can be used to reduce dimensionality to $\sim 10$ or $\sim 50$.
@@
The phylogeny of languages might also not be the aspect most correlated with the distances we see. For example geographical and structural similarities as studied in Bjerva et al. (2019) could lead to closer correlations.
Our results demonstrate that TDA invariants capture properties of languages, but what aspects of language are they keying on? What do tunnels and voids in embeddings tell us about languages?
@@ Appendix A Supplemental Material
A.1 Vietoris-Rips complex
To compute persistent homology, the continuous spaces of growing balls around the points in the cloud need to be replaced by a discrete structure easily represented in a computer. Concretely, by a simplicial complex, a (hyper)graph-like structure. A simplicial complex consists of a set of vertices, and a set of faces. Each face is a subset of vertices, and for each face, a simplicial complex has to also include all its subsets. A filtration of a simplicial complex is a function that assigns a real number to each face, and is monotone with respect to inclusion, that is, a subset of a face has to have smaller or equal value. A collection of all faces with value at most $r\geq 0$ is then a simplicial complex. This yields a chain of growing simplicial complexes.
For completeness we provide the formal definition of a ‘feature’ within a VR complex. A boundary $\partial\sigma$ of a face $\sigma\in\mathrm{VR}_{r}({D})$ is the collection of subsets $\sigma\setminus\{i\}$ for each $i\in\sigma$. An addition operation is defined on boundaries as the symmetric difference of the collections. A boundary of a collection of faces is then the sum of their boundaries. For example, the boundary of two triangles sharing an edge are the four edges on the outside.
There is a bar $[r_{b},r_{d})$ in the $p$-dimensional persistence diagram of $\mathrm{VR}_{\bullet}({D})$ iff there exist two collections of faces, $\mathcal{B}\subseteq\mathrm{VR}_{r_{b}}({D})$ and $\mathcal{D}\subseteq\mathrm{VR}_{r_{d}}({D})$, with the following properties. First, the boundary $\partial\mathcal{B}$ is empty, and one of the faces in $\mathcal{B}$ appears for the first time in $\mathrm{VR}_{r_{b}}({D})$. Second, the boundary $\partial\mathcal{D}$ is $\mathcal{B}$, and one of the faces in $\mathcal{D}$ appears for the first time in $\mathrm{VR}_{r_{d}}({D})$. For example, a $1$-dimensional bar means that an edge appears in $r_{b}$ which completes a loop, and then this loop is filled by a collection of triangles, the last of which appears in $r_{d}$.
A.2 Preserving the Ethnologue Tree Topology
@@
Thus we investigated an optimization procedure to construct TDA-based trees consistent with the reference tree topology. Instead of constructing a new tree, we try to find the best assignment of labels (languages) to the leaves based on the given distance matrix. A tree always yields a distance matrix between the leaves—the distance of two leaves is the length of the unique path connecting them. We set up an optimization process to find the label assignment leading to the tree distance matrix that best correlates with our given distance matrix. To compare distance matrices, we use the Pearson correlation coefficient. Therefore, if $E$ is the tree distance matrix of the Ethnologue tree, and $D$ is the distance matrix computed from the pipeline as described above, we want to find a permutation matrix $P$ that maximizes the Pearson correlation coefficient between $P^{T}EP$ and $D$.
@@
A.3 Tree Similarity Metrics